What is a sentence for the word latent?
The forensic scientist tested the crime scene for latent fingerprints.
What are uses of ultrasound in day to day life?
Ultrasound are usedful in diagnosing medical problems caused to human beings.
What is the probability of rolling a 6 on a die two consecutive times?
Two dice can come up in (6 x 6) = 36 possible ways.
There are six ways to roll a 6:
1, 5
2, 4
3a, 3b
3b, 3a
4, 2
5, 1
The probability of rolling a six on the first toss is (6/36) = 1/6
The probability of rolling a six on the second toss is (6/36) = 1/6
The probability of both is (1/6) x (1/6) = 1/36 = 2.78% (rounded)
What does standard deviation tell about distribution?
How widely spread out, or tightly concentrated about the mean it is.
What formula is used to find the theoretical probability of an event?
Expected successes= Theoretical Probability · Trials
P(event) = Number of possible out comes divided by total number of possible
Will robots exist in 25 years?
Yes. Robots exist now, and there is no reason to believe that they will stop existing.
If, however, you meant to ask "Will intelligent robots like those in some science fiction novels or movies, like Star Wars, exist in 25 years?", that is not known, but it seem unlikely, given the immense complexity of the human brain.
What is mean by standard recipes?
The standard recipes refer to the ideal recipes that is used in the normal situation.
What is the probability that a random card picked up from a deck is a 10 or a spade?
There are 4 tens and 13 spades but one of these is the ten of spades.
So there are 4 + 13 - 1 = 16 cards which lead to "success".
Therefore the probability is 16/52 = 4/13
What is an outcome in a probability experiment?
It is the result of the experiment. It is the value of the observation.
How can you make theoretical probability more accurate?
By ensuring your model is as good as it can be. Make sure that any assumptions that you make for your model are justified and, if necessary, properly reflected in the model.
What are the chances of having a girl or boy?
1. The Chances of having a baby boy and a baby girl is 50/50 because the girls have XX Chromosomes and Boys have XY Chromosomes it's a 50% chance that it sprays some liquid to get the Y Chromosomes so either way its 50/50 chance
What is the probability that both numbers on the dice are less than 4?
Any normal die contains three numbers that are less than 4 so the answer must be 1.
What is the probability that a cube will land on 4. A coin will land on heads?
Prob(Cube = 4) = 1/6
Prob(Coin = H) = 1/2.
Prob(Cube = 4) = 1/6
Prob(Coin = H) = 1/2.
Prob(Cube = 4) = 1/6
Prob(Coin = H) = 1/2.
Prob(Cube = 4) = 1/6
Prob(Coin = H) = 1/2.
What is the montry hall problem?
Perhaps you mean "Monty Hall" problem. If so, see the related question:
What is the Monty Hall Paradox?
mean
What are the odds of picking a king or a red card?
I calculated this assuming you do not have jokers in the deck.
Getting a red card is easy because that is half the deck. 50% chance or 1/2 chance of getting a red card on the next drop. If its a heart or a diamond (specifically) you are looking for its a 25% or 1/4 chance. (also a 75% or 3/4 chance it WON'T be the suit you are looking for)
Next is the Kings. 52 (cards in a deck) divided by 100 is 0.52. 0.52 multiplied by 4 (four kings in a normal deck) is 2.08. So at all times a 2.08% chance a king will drop. Keep in mind the opposite is true, that there is like 97.92% chance that a king wont drop. The "odds" of this are approximately 2/100 or 1/50. Someone please feel free to correct me and my math.
In mathematical terms what is replacement?
Replacement mean we put "it" back before the next step. For example, in drawing cards, replacement means we put the card back before we draw another one. So given an arbitrary collection of objects we want to know what the probability of an event is with replacement. This is equivalent to saying what is the probability of an event when we remove an item from the collections, replace it, mix up the collection, and remove an item again. We can continue doing this infinite number of times.
Implementation on MIS in a distribution firm?
thinking what do you need to know as a chief manager of distribution firm let say this Instrumentation products firm
1. No of employees in your firm, their capabilities, make heirarchy among the employees
2. the area which are your firm is in
3. the number of dealers you have for your firm
4. The end customers in different regions
5. Sources to know about the movement of products
If I were just a chilef manager and do not have the subjective knowledge, I would have proceed in this way.
If a coin is flip 9 times and get all H What is the probability of an H on next flip?
The probability of the coin landing "head" side up is 50/50, meaning it could land "head" side up or "tail" side up. The odds of any single coin flip are always the same, no matter what happened on the previous tosses -- provided the coin is not a "double-head" (or "double-tail") "trick" coin
What is Construction of a sampling distribution from scratch is feasible when the population is?
Real
The area under the normal curve is greatest in which scenario?
The area under the normal curve is ALWAYS 1.